OpenStudy (anonymous):

what is the derivative of te^-t/2

OpenStudy (australopithecus):

te^(-t/2) the derivative of f(x) = e^u f'(x) = (e^u)u' Product rule is d(x) = f(x)g(x) d'(x) = f'(x)g(x) + g'(x)f(x) split the problem in to smaller functions. so find the derivative of t then find the derivative of e^(-t/2) then apply product rule

OpenStudy (australopithecus):

pretend f(x) = e^(-t/2) and g(x) = t then just take the derivative of these two functions and sub that into the product rule